OpenStudy (anonymous):

One times a number x, plus 2 times a number y, equals 9. Negative one times x minus 3 times y equals -14. What are the numbers?

OpenStudy (anonymous):

A. x=7, y=12 B. x=-1, y=5 C. X=5, y=1 D. x=-11, y=24

OpenStudy (anonymous):

Help me @mathmale

OpenStudy (anonymous):

I'll give medal

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

@texaschic101

OpenStudy (kkutie7):

Lets start by setting this up =) first sentence: \[(1*x)-(2*y)=9\rightarrow x+2y=9\] \[(-1*x-(3*y)=-14\rightarrow -x-3y=-14\]

OpenStudy (kkutie7):

now lets solve for y the first is \[x+2y=9\rightarrow 2y=9-x\rightarrow y=\frac{9-x}{2}\] the second is \[-x-3y=-14\rightarrow -3y=-14+x\rightarrow 3y=14-x\rightarrow y=\frac{14-x}{3}\]

OpenStudy (kkutie7):

\[\frac{14-x}{3}=\frac{9-x}{2}\rightarrow 2(14-x)=3(9-x)\rightarrow\] \[28-2x=27-3x\rightarrow 28-27=-3x+2x\rightarrow 1=-x\rightarrow x=-1\]

OpenStudy (anonymous):

Two times a number, x, plus 3 times a number, y, equals 50. Four times x minus 2 times y equals 4. What are the numbers? x = 7, y = 12 x = 19, y = 4 x = 10, y = 18 x = -11, y = 24

OpenStudy (kkutie7):

\[y=\frac{9-(-1)}{2}\rightarrow \frac{10}{2}\rightarrow y=5\] \[y=\{14-(-1)}{3}\rightarrow \frac{15}{3}\rightarrow y=5\]

OpenStudy (anonymous):

Thank you sorry could you help me on this last one to its my last one out of 80 questions

OpenStudy (anonymous):

It's 2:00 a.m 😞

OpenStudy (kkutie7):

quick answer or you want a walk through?

OpenStudy (anonymous):

Quick answer

OpenStudy (kkutie7):

x=11/2 and y=9

OpenStudy (kkutie7):

I didnt look at you answers so don't know if I'm remotely correct I'm tired.

OpenStudy (anonymous):

Ok thank you anyway for Your time so helpful

OpenStudy (kkutie7):

i try